P, = 500 lb 8 ft 2 ft 2 ft BFind the forces on each member
Answer #1

Expressing the force is each member is vector form. A(3,-4.0) D(0.0,8) C(-2,2,0) B(3,2,0) AD=-31 +4j+8k BD=-31-2; +8k CD=-21-For equilibrium at joint D we have sum of all the forces(EF)=0. Pz + Fod+F8D + FcD =0 -FO(-0.31791 +0.4237; +0.8479k) - FBD(-Considering the force body diagram of joint B Z Го? FBC Feet g Br Br For equilibrium EF=0 Fax +Fy+Fpc +B=0 (BC BA BA + FOD BD